# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v2

@@ Appendix B FinSABER Strategies Base
B.1 Timing-based Strategies
@@
This category includes FinMem (Yu et al., 2023) and FinRobot (Yang et al., 2024). We acknowledge other works, such as FinAgent (Zhang et al., 2024b), FinCon (Yu et al., 2024) but they are not (yet) open-source, which prevents us from generating backtesting results.
@@
RL-Based Strategies
Additionally, we incorporate RL-based strategies to evaluate their performance in investing environments. Specifically, we implement Advantage Actor-Critic (A2C), Proximal Policy Optimization (PPO), Deep Deterministic Policy Gradient (DDPG), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), leveraging the FinRL framework (Liu et al., 2022, 2024). These RL algorithms are widely used in financial markets due to their ability to learn optimal investing policies from historical data and adapt to evolving market conditions. A2C and PPO, as policy-gradient methods, optimise investing decisions while ensuring stability, whereas DDPG, TD3, and SAC extend this capability to continuous action spaces, rendering them well-suited for asset allocation and portfolio optimisation tasks.
B.2 Selection-based Strategies
(1) Random Five, which involves randomly selecting five stocks from historical S&P 500 constituents available at each rolling evaluation period’s start date, (2) Momentum Factor Selection, which selects stocks based solely on past returns at each period’s start date, following a momentum factor-based strategy (Muller and and, 2010), and (3) Volatility Effect Selection, which selects stocks with the lowest historical return volatility, capturing the empirically observed inverse relationship between volatility and risk-adjusted performance (Blitz and Vliet, 2007).
@@ Appendix C Evaluation Metrics
We group evaluation metrics into three categories, each targeting a distinct aspect of strategy performance:
C.1 Return Metrics
Annualised Return (AR).
Measures the geometric average yearly return: $R_{\textrm{annual}}=(1+C)^{\frac{252}{T}}-1$ where $C=\prod_{t=1}^{T}(1+R_{t})-1$.
Cumulative Return (CR).
Total portfolio return over the test period: $C=\prod_{t=1}^{T}(1+S_{t}\cdot R_{t})-1$, where $S_{t}$ is the position and $R_{t}$ is the market return.
C.2 Risk Metrics
Annualised Volatility (AV).
Volatility of daily returns, scaled annually: $\sigma_{\textrm{annual}}=\sigma_{\textrm{daily}}\times\sqrt{252}$.
Maximum Drawdown (MDD).
Largest drop from peak portfolio value: $\textrm{MDD}=-\max_{t\in[1,T]}\left(\frac{\textrm{Peak}_{t}-V_{t}}{\textrm{Peak}_{t}}\right)$.
C.3 Risk-adjusted Performance Metrics
Sharpe Ratio (SPR).
Return per unit of total volatility: $\textrm{SPR}=\frac{\overline{R_{t}}-\frac{r_{f}}{252}}{\sigma_{\textrm{daily}}}\times\sqrt{252}$.
Sortino Ratio (STR).
Return per unit of downside deviation: $\textrm{STR}=\frac{\overline{R_{t}}-\frac{r_{f}}{252}}{\sigma_{\textrm{down}}}\times\sqrt{252}$.
Where $\overline{R_{t}}$ is the average daily return, $r_{f}$ is the annual risk-free rate, and $\sigma_{\textrm{down}}$ is the standard deviation of negative returns.
@@ Appendix D Extra Results on Selective Symbols
Tables 2 and 5 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (Yu et al., 2023) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from Yu et al. (2024), while the traditional rule-based results presented are based on our implementations.
@@
Rule Based Turn of The Month 0.928 27.850 -11.642 0.123 3.487 -14.892 0.874 31.344 -17.995 0.407 7.744 -11.955
@@
Rule Based Turn of The Month 0.098 3.337 -12.498 0.343 7.188 -13.519 0.987 18.942 -10.641 -0.020 -8.999 -33.895
@@
Table 5: Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in Yu et al. (2024), evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.
@@ Appendix E Technical Details
RL total_timesteps=50000
@@
Table 6: Default parameter settings for benchmark strategies.
Appendix F LLM Strategies Cost Analysis
@@
FinAgent was roughly 6 times more expensive than FinMem in our tests. Importantly, these figures only account for LLM generation costs (i.e., chat/completions endpoints), and do not include the cost of generating embeddings (e.g., via text-embedding-ada-002), which would further increase the total budget.

### v3

@@ Appendix B FinSABER Strategies Base
B.1. Timing-based Strategies
@@
This category includes FinMem (Yu et al., 2023) and FinRobot (Yang et al., 2024). We acknowledge other works, such as FinCon (Yu et al., 2024) and MarketSenseAI (Fatouros et al., 2025), but they are not (yet) open-source, which prevents us from generating backtesting results.
@@
RL-Based Strategies.
We also implement widely used RL algorithms for financial markets, including Advantage Actor-Critic (A2C), Proximal Policy Optimisation (PPO), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), utilising the FinRL framework (Liu et al., 2022, 2024). Each agent learns investing policies by interacting with a simulated trading environment based on the OpenAI Gym API, using real historical market data.
  • Environment design. The trading process is modelled as a Markov Decision Process (MDP), where the agent observes a market state, selects an action, and receives a reward at each step. The environment simulates live stock markets and follows the time-driven simulation principle.
  • State space. At each timestep, the agent receives a state vector capturing account balance, current shares held for each asset, stock prices (close, open, high, low), trading volumes, and technical indicators (e.g., MACD, RSI). The feature set is configurable and can be aggregated at daily, hourly, or minute frequency.
  • Action space. The action space allows buying, holding, or selling any subset of assets. Typically, actions are discrete values representing the number of shares to transact per asset (e.g., $a\in\{-k,\ldots,-1,0,1,\ldots,k\}$ for each stock), where negative, zero, and positive values denote selling, holding, and buying, respectively.
  • Reward function. The primary reward is the change in portfolio value after each trading action, i.e., $r(s,a,s^{\prime})=v^{\prime}-v$, where $v$ and $v^{\prime}$ are the portfolio values at states $s$ and $s^{\prime}$.
B.2. Selection-based Strategies
This section details the implementation of the primary selection strategies used in our composite backtesting framework. Each selector operates on the historical S&P 500 constituents available at the start of a given rolling-window period to produce a list of tickers for the timing-based strategies.
Random Five
This strategy serves as a simple baseline for performance comparison. At the beginning of each evaluation period, it selects five stocks at random, without replacement, from the list of all available historical S&P 500 constituents for that period.
Momentum Factor
Following the well-documented momentum factor (Muller and and, 2010), this strategy selects the stocks with the highest recent price appreciation. For each candidate stock, we calculate a momentum score based on its historical price data. Specifically, the score is the percentage return over a “momentum period” (e.g., 100 trading days), but we exclude the most recent “skip period” (e.g., 21 trading days) from the calculation. This practice is common in momentum strategies to avoid the “short-term reversal” effect (Carhart, 1997). The score for a given stock is calculated as shown in Equation 1. $t$ is the selection date. All candidate stocks are then ranked in descending order by this score, and the top-$k$ stocks (e.g., $k=5$) are selected.
(1) $\text{Momentum Score}=\frac{\text{Price}_{t-\text{skip\_period}}}{\text{Price}_{t-\text{momentum\_period}}}-1$
Volatility Effect
This strategy is based on the “volatility effect” anomaly, where low-volatility stocks have been empirically shown to generate higher risk-adjusted returns (Blitz and Vliet, 2007). For each candidate stock, we measure its historical volatility over a recent “look-back period” (e.g., 21 trading days). The volatility is calculated as the standard deviation of its weekly log returns within this period. We use weekly returns ($\ln(P_{t}/P_{t-5})$) rather than daily returns to smooth out daily noise. Candidate stocks are then ranked in ascending order by their calculated volatility, and the top-$k$ stocks with the lowest volatility are selected for the portfolio.
FinCon Selection Agent.
Unlike the single-factor methods above, the FinCon Selection Agent (Yu et al., 2024) aims to construct a diversified portfolio by explicitly considering both performance and inter-stock correlation. Its selection process is more sophisticated:
  1. (1)
    Metric Calculation: For all candidate stocks over a “look-back years” period (e.g., 2 years), the agent calculates daily returns to derive a full correlation matrix and a suite of performance metrics for each stock, including the Sharpe ratio.
  2. (2)
    Primary Selection: The agent first ranks each stock using a combined score that balances risk-adjusted return (Sharpe ratio) and its potential for diversification (low average correlation with all other stocks, $\overline{\rho}$). The score is calculated as:
    $\text{Score}=\text{Sharpe Ratio}\times(1-\overline{\rho})$
    The top-$k$ stocks based on this score form the initial portfolio.
  3. (3)
    Diversification Check & Fallback: The agent then assesses the average correlation within the selected $k$-stock portfolio. If this internal correlation is above a predefined threshold (e.g., 0.7), it indicates poor diversification. In this case, the agent discards the initial selection and triggers a fallback algorithm. This second algorithm uses a greedy, diversification-first approach: it starts with the single stock with the highest Sharpe ratio and then iteratively adds the available stock that has the lowest average correlation to the already-selected members until a $k$-stock portfolio is formed.
This two-stage process ensures that the final selection prioritises a balance of individual stock quality and overall portfolio diversification.
@@ Appendix C Evaluation Metrics
We group evaluation metrics into three categories, each targeting a distinct aspect of strategy performance. In the following definitions, $T$ represents the total number of trading days, and $R_{t}$ is the portfolio’s return on day $t$.
C.1. Return Metrics
Annualised Return (AR)
Measures the geometric average return of the portfolio on a yearly basis. It is calculated from the total cumulative return $C$ as:
(2) $R_{\textrm{annual}}=(1+C)^{\frac{252}{T}}-1$
where 252 is the approximate number of trading days in a year.
Cumulative Return (CR)
Measures the total return of the portfolio over the entire test period. It is calculated as:
(3) $C=\prod_{t=1}^{T}(1+S_{t}\cdot R_{m,t})-1$
where $S_{t}$ is the position taken by the strategy on day $t$ (+1 for long, 0 for neutral) and $R_{m,t}$ is the market return of the asset on day $t$.
C.2. Risk Metrics
Annualised Volatility (AV)
Measures the standard deviation of the portfolio’s returns, scaled to a yearly figure. It is defined as:
(4) $\sigma_{\textrm{annual}}=\sigma_{\textrm{daily}}\times\sqrt{252}$
where $\sigma_{\textrm{daily}}$ is the standard deviation of the portfolio’s daily returns, $R_{t}$.
Maximum Drawdown (MDD)
Measures the largest peak-to-trough decline in portfolio value, representing the worst-case loss from a previous high. It is defined as:
(5) $\textrm{MDD}=\max_{t\in[1,T]}\left(\frac{P_{t}-V_{t}}{P_{t}}\right)$
where $V_{t}$ is the portfolio value on day $t$, and $P_{t}$ is the peak portfolio value recorded up to day $t$ ($P_{t}=\max_{i\in[1,t]}V_{i}$).
C.3. Risk-adjusted Performance Metrics
Sharpe Ratio (SPR)
Measures the excess return of the portfolio per unit of its total volatility. It is calculated as:
(6) $\textrm{SPR}=\frac{\overline{R_{t}}-R_{f,\text{daily}}}{\sigma_{\textrm{daily}}}\times\sqrt{252}$
Sortino Ratio (STR)
Similar to the Sharpe ratio, but it only penalises for downside volatility, measuring the excess return per unit of downside risk. It is defined as:
(7) $\textrm{STR}=\frac{\overline{R_{t}}-R_{f,\text{daily}}}{\sigma_{\textrm{downside}}}\times\sqrt{252}$
where $\overline{R_{t}}$ is the average daily portfolio return, $R_{f,\text{daily}}$ is the daily risk-free rate (i.e., the annual rate divided by 252), and $\sigma_{\textrm{downside}}$ is the standard deviation of only the negative daily returns.
@@ Appendix D Extra Results on Selective Symbols
Tables 2 and 7 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (Yu et al., 2023) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from Yu et al. (2024), while the traditional rule-based results presented are based on our implementations.
@@
Rule
Based
Turn of The Month 0.928 27.850 -11.642 0.123 3.487 -14.892 0.874 31.344 -17.995 0.407 7.744 -11.955
@@
Rule
Based
Turn of The Month 0.098 3.337 -12.498 0.343 7.188 -13.519 0.987 18.942 -10.641 -0.020 -8.999 -33.895
@@
Table 7. Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in Yu et al. (2024), evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.
@@ Appendix E Technical Details
RL-A2C learning_rate=1e-5, ent_coef=0.1, vf_coef=0.5, max_grad_norm=0.5, gae_lambda=0.95, gamma=0.99
RL-PPO batch_size=64, learning_rate=2.5e-4, ent_coef=0.1, clip_range=0.2, gae_lambda=0.95, gamma=0.99
RL-SAC learning_rate=2e-2, buffer_size=1000000, batch_size=256, learning_starts=100, ent_coef=0.1, tau=0.005, gamma=0.99, action_noise=”normal”
RL-TD3 learning_rate=3e-2, buffer_size=1000000, tau=0.005, gamma=0.99, policy_delay=2, target_policy_noise=0.5, target_noise_clip=0.5, action_noise=”normal”
@@
Table 8. Default parameter settings for benchmark strategies.
Appendix F Comparative Drawdown Analysis via Underwater Plots
This appendix provides a visual analysis of strategy risk profiles through underwater plots. An underwater plot visualises the drawdown of a portfolio over time, offering an intuitive way to assess the depth, duration, and frequency of its losses.
The plots are derived by calculating the percentage loss of a portfolio’s equity curve from its running maximum value (its previous peak). At any given point in time, the drawdown $D_{t}$ is calculated as:
(8) $D_{t}=\frac{\text{Current Value}_{t}-\text{Previous Peak}_{t}}{\text{Previous Peak}_{t}}$
A value of 0% indicates the portfolio is at a new all-time high, while a negative value shows how far it is “underwater”. When interpreting the plots, two key features should be considered:
  • Depth: The magnitude of the drawdown, indicated by how low the line drops on the y-axis. Deeper drawdowns represent larger losses and greater risk.
  • Duration: The length of time the line stays below the 0% axis. Longer durations represent slower recoveries and more prolonged periods of underperformance for the investor.
A superior strategy will exhibit shallower and briefer drawdowns compared to its benchmark.
Figure 3. Comparative underwater plots for the FinMem (blue) and FinAgent (red) strategies against the Buy and Hold (SPX) benchmark across individual stocks selected in the Composite setup. The plots are grouped by the market regime of the period shown: bull markets (top two rows), bear market (third row), and sideways markets (bottom two rows).
The visual case studies shown in Figure 3 complement the aggregated quantitative results in the main paper, offering a granular perspective on the agents’ behavioural patterns under different market conditions.
Bull Markets.
The top two rows of the figure display strategy performance during bull market years, revealing a stark divergence in the agents’ approaches. The FinAgent strategy (red) sometimes exhibits an overly conservative posture, as seen in KO (2019-2020) and APA (2023-2024). Its drawdowns are shallower than the benchmark’s, or it may not trigger any trading activities. While this appears safe, it visually confirms the low beta values from our quantitative analysis and indicates a missed opportunity to capitalise on market gains. However, this risk-averse behaviour is fragile; in the case of ULTA (2019-2020), FinAgent experiences a catastrophic drawdown, revealing its risk model to be unreliable and poorly calibrated.
In contrast, the FinMem strategy (blue) consistently fails to manage single-stock volatility. In most bull-market cases (DOV, WHR, DXC), its drawdowns are significantly deeper and more prolonged than FinAgent’s. This demonstrates an inability to handle the inherent risk of the underlying asset, leading to the significant underperformance identified in the main paper.
Bear Markets.
The third row, depicting the 2008 Global Financial Crisis, provides the most critical insight into the agents’ flaws. While a single stock is expected to be more volatile than the index during a crash, the LLM strategies, particularly FinMem, catastrophically amplify this downside risk. For DE, the FinMem strategy’s drawdown approaches -75%, a far more severe loss than the SPX benchmark’s -50%. Rather than providing any form of risk mitigation, the agents appear to make pro-cyclical decisions that accelerate losses. The FinAgent strategy, true to its more conservative nature, often mitigates some of these losses relative to FinMem, yet it still fails to generate a positive outcome. For instance, while its drawdown for CNX is shallower than FinMem’s, it remains severe and prolonged. This relative outperformance is insufficient and aligns with our market regime analysis (§7), which finds that both agents are poorly calibrated for bear markets and ultimately succumb to losses (Cont, 2001).
Sideways Markets.
The final two rows illustrate performance in sideways, where the primary challenge is managing idiosyncratic stock risk without a clear market tailwind. Generally (but not consistently), the FinAgent strategy (red) exhibits shallower and less severe drawdowns than FinMem (blue), as seen in cases like NCR (2004-2005), FDO (2009-2010), and SYY (2016-2017). However, FinAgent’s conservative nature can also lead to periods of complete inactivity where no trades are triggered (observed before in bull market and bear market), causing it to miss minor recovery opportunities that the benchmark captures, as seen in FDO (2009-2010).
In summary, these visual case studies reinforce the quantitative conclusions in §6.3 and §7. LLM agents are poorly calibrated to distinct market regimes, behaving too timidly in uptrends and too recklessly in downturns, ultimately failing to provide the adaptive risk management necessary for consistent performance.
Appendix G LLM Strategies Cost Analysis
@@
FinAgent was roughly 6 times more expensive than FinMem in our tests. Importantly, these figures only account for LLM generation costs (i.e., chat/completions endpoints), and do not include the cost of generating embeddings (e.g., via text-embedding-ada-002111111https://platform.openai.com/docs/models/text-embedding-ada-002), which would further increase the total budget.